reviews the approach that is necessary for solving problems such as this one. A baseball (m = 145 g) approaches a bat horizontally at a speed of 42.1 m/s (94.2 mi/h) and is hit straight back at a speed of 46.5 m/s (104 mi/h). If the ball is in contact with the bat for a time of 1.30 ms, what is the average force exerted on the ball by the bat? Neglect the weight of the bat, since it is so much less than the force of the bat. Choose the direction of the incoming ball as the positive direction.
Answer #1

F = dp/dt OR change in momentum / ( change in time)

=> F = - 0.145*( 46.5+42.1)/(1.30*10^-3) = - 9882.3 N